# Forging Tokens for Improved Storage-efficient Training

### v1

@@ Appendix C More qualitative examples
We demonstrate the effectiveness of the proposed TokenAdapt and ColorAdapt by decoding the augmented tokens into images using the ViT-VQGAN decoder [56]. In Figure A, we compare the direct application of pixel-based data augmentations to token embedding with augmentations using our TokenAdapt module. As mentioned in Section 3.1, direct application of hFlip or affine transformations causes disruption in both the diagonal line and the object’s silhouette due to spatial information collapse. Additionally, data augmentations related to interpolation (e.g., Resize or Mixup) can result in undesired artifacts. Mixup, in particular, occasionally causes significant disruption of objects, leading to substantial performance degradation as shown in Table 1. The occurrence of these unexpected artifacts makes it difficult to use the data augmentations, which are widely used in existing pixel image domains. In contrast, The tokens that are augmented using our TokenAdapt module exhibit more reasonable results, mitigating the image degradation caused by the direct application of pixel-based augmentations to tokens. Consequently, TokenAdapt consistently improves model performance across various scenarios by effectively leveraging pixel-based data augmentation in a token domain.
@@
In Figure B, we present the images decoded from color-augmented tokens by various existing color-based data augmentations. For pixel-based color augmentation, we employ brightness and contrast, which are widely adopted for pixel-based vision model training. Specifically, we used brightness and contrast functions following the implementation [38]. For token-based color augmentation, we employ Emb-Noise [36] with the same optimized hyperparameters used in token-based vision model training [36]. Figure B illustrates the impact of the proposed ColorAdapt on the visual characteristics of augmented tokens. Unlike the existing color-based augmentations, our ColorAdapt effectively preserves object structure while introducing significant color variations. Recognizing the importance of maintaining object structure in diverse vision tasks (e.g., fine-grained classification, semantic segmentation), our ColorAdapt opens up new possibilities for training more robust and adaptable vision models.

### v3

@@ Appendix 0.B More Qualitative Examples
We demonstrate the effectiveness of the proposed TokenAdapt and ColorAdapt by decoding the augmented tokens into images using the ViT-VQGAN decoder [59]. In Figure C, we compare the direct application of pixel-based data augmentations to token embedding with augmentations using our TokenAdapt module. As mentioned in Section 3.3, direct application of hFlip or affine transformations leads to disruptions in both the diagonal line and the silhouette of the object due to spatial information collapse. In addition, data augmentations related to interpolation (e.g., Resize or Mixup) can result in undesired artifacts. Mixup, in particular, occasionally causes significant disruption of objects, leading to substantial performance degradation as shown in Figure 3(b).
The occurrence of these unexpected artifacts makes it difficult to use the data augmentations that are widely used in existing pixel image domains. Moreover, in tasks such as MTM and semantic segmentation, where pixel-level details are essential, such artifacts significantly affect learning stability and model performance. In contrast, The tokens that are augmented using our TokenAdapt module exhibit more reasonable results, mitigating the image degradation caused by the direct application of pixel-based augmentations to tokens. Consequently, TokenAdapt consistently improves model performance across various scenarios by effectively leveraging pixel-based data augmentation in a token domain. Notably, as shown in Table 2, we observe that SeiT++ shows more significant performance gains with MTM, highlighting the importance of minimizing undesired artifacts to overcome the challenges of data augmentation in the tasks where pixel-level information is important.
0.B.0.2 ColorAdapt.
@@
In Figure D, we present the images decoded from color-augmented tokens by various existing color-based data augmentations. We employ brightness and contrast for pixel-based color augmentation, which are widely adopted for pixel-based vision model training. Specifically, we used brightness and contrast functions following the implementation [40]. For token-based color augmentation, we employ Emb-Noise [38] with the same optimized hyperparameters used in token-based vision model training [38]. Figure D illustrates the impact of the proposed ColorAdapt on the visual characteristics of augmented tokens. Unlike the existing color-based augmentations, our ColorAdapt effectively preserves object structure while introducing significant color variations. Recognizing the importance of maintaining object structure in diverse vision tasks (e.g., fine-grained classification, semantic segmentation), our ColorAdapt opens up new possibilities for training more robust and adaptable vision models.
Appendix 0.C Implementation Details
0.C.0.1 Masked Token Modeling.
For masked token modeling (MTM), we follow the training recipe from MAGE [32], adjusting the pre-training epochs and the masking ratio. We pre-train the ViT-B model [17] for 400 epochs with a batch size of 4096 and fine-tune the model for 100 epochs with a batch size of 1024. During pre-training, we randomly mask out certain input tokens with a variable masking ratio ranging from 0.4 to 1. The base learning rate is 0.00015 and 0.001 for pre-training and fine-tuning, respectively. In MTM, we replace the Conv Stem-Adapter module from SeiT [38] with Conv as the patch embedding layer for ViT models. This adjustment is made because Conv creates overlapping input patches, which hinders representation learning based on masked token modeling. Regarding data augmentation, recognizing the crucial role of pixel-level information in MTM, we apply geometric pixel-data augmentations using only the proposed TokenAdapt module, enhancing our training paradigm’s effectiveness.
0.C.0.2 Token-based Image Classification.
For token ImageNet-1k training, we follow the training recipe from SeiT [38], adjusting only the warm-up epochs for stable convergence. We conducted token-based image classification on the ViT-B/16 model [17, 50] and used a learning rate of 0.0015 with cosine scheduling and a weight decay of 0.1. The model was trained for 300 epochs with a batch size of 1024. Regarding data augmentation, SeiT [38] incorporates Token-RRC, Token-EDA, Token-CutMix, and Emb-Noise. As a default, we integrate these token-based data augmentation methods and further enhance token diversity by applying additional augmentations using our TokenAdapt and ColorAdapt. Specifically, we apply pixel-based data augmentations (e.g., RRC, hFlip, affine transformations, Mixup [61], and CutMix [60]) to tokens using our TokenAdapt module with a probability of 0.5. We adopt the hyperparameters for data augmentation proposed in DeiT [50]. For token-based fine-grained classification, following SeiT, we use DeiT’s training recipe. When the number of data points decreased in all experiments, we adjust the number of total training iterations to ensure a fair comparison.
0.C.0.3 Token-based Semantic Segmentation.
For the tokenized ADE-20k dataset preparation, we initially resize the entire ADE-20k dataset to . Following the procedure described in SeiT, we use the ImageNet-1k-trained ViT-VQGAN tokenizer [59] to extract tokens from the resized images. Token-based semantic segmentation on ADE-20k follows the training recipe of mmsegmentation [9], using the DeiT-B/16 model with UperNet [56]. The training involves a learning rate of 6e-5 with polynomial scheduling and a weight decay of 0.01. The model was trained for 80k iterations with a batch size of 16. Regarding data augmentation, we do not employ mixup-based augmentations in token-based semantic segmentation. This behavior is consistent with recent studies [57, 48, 6] that exclude mixup-based augmentation in their training recipes. Furthermore, similar to MTM, we observe that token-level Random Resized Crop (RRC) does not improve performance due to the importance of maintaining structural integrity in pixel-level classification tasks. Thus, we apply geometric pixel-based data augmentations only using our TokenAdapt module.